OpenStudy (anonymous):

How do you find dy/dx in terms of x and y if xln(y) + y^(5) = 9ln(x) ?

OpenStudy (anonymous):

Hey there! I take it you are probably in Calc BC? The times! The first thing that is very simple to do is get all terms on one side. so ys on the left and xs on the right.

OpenStudy (anonymous):

I hope this helps!

OpenStudy (anonymous):

I'm actually in Calc 1, haha. I got \[\frac{ x }{ y } + \ln(y) - \frac{ 9 }{ x } = -5y ^{4}\] But I'm stuck there :/

OpenStudy (anonymous):

Actually I don't even know if that's right. I just took the derivative of both sides?

OpenStudy (anonymous):

Fancy Stuff. You still got a few things to move over. Just from the view of things, you got a lonely ln(y) thats wants to be with a friend

OpenStudy (anonymous):

What can I do with \[-5y ^{4} - \ln(y)\] ?

OpenStudy (anonymous):

Nevermind, I got it. Thanks! :)

OpenStudy (anonymous):

This would be a separation of variables.

OpenStudy (anonymous):

Seperate the terms to have all of Y on one side, and all of X on the other.

OpenStudy (noelgreco):

Use implicit differentiation.